# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 8.6. An example of formerly open complexity
(recall that $((x_{1},y_{1}),(x_{2},y_{2}))$ might have been added to $R^{{\mathfrak{M}}}$ so that ${\mathfrak{M}}$ satisfies the sentence in (21)). The last option is false by the assumption and by (19), ${\mathfrak{B}}\models\neg S(x_{2})\wedge\neg S(y_{2})$, and hence neither (B) nor (D) holds. Therefore, one of the conditions (22), (A), or (C) holds for $((x_{1},y_{1}),(x_{2},y_{2}))$. Similarly, we obtain that one of the conditions (22) or (B) holds for $((x_{2},y_{2}),(x_{1},y_{1}))$, since ${\mathfrak{M}}\models R((x_{2},y_{2}),(x_{1},y_{1}))$ (to exclude (D) we use the assumption that ${\mathfrak{B}}\models S(x_{1})$ and hence $x_{1}\neq x_{2}$). This yields six cases and in each of them we must have that ${\mathfrak{B}}\models R(x_{1},x_{2})\wedge R(x_{2},x_{1})$ or ${\mathfrak{B}}\models S(y_{1})\wedge R(y_{1},y_{2})\wedge R(y_{2},y_{1})$. Since ${\mathfrak{B}}\models S(x_{1})\wedge R(x_{2},x_{2})\wedge R(y_{2},y_{2})$, this contradicts ${\mathfrak{B}}\models\neg\mu$. Since $(x_{1},y_{1}),(x_{2},y_{2})\in M$ were chosen arbitrarily, this shows that ${\mathfrak{M}}\models\neg\mu$. The argument for ${\mathfrak{N}}$ is similar.
Claim 5.
There is an embedding $f$ of ${\mathfrak{M}}$ into ${\mathfrak{B}}$ and an embedding $g$ of ${\mathfrak{N}}$ into ${\mathfrak{B}}$. We show the claim for ${\mathfrak{M}}$; the proof for ${\mathfrak{N}}$ is analogous. By (Bodirsky, 2021, Lemma 4.1.7), it is enough to show that every finite substructure of ${\mathfrak{M}}$ embeds into ${\mathfrak{B}}$. By the definition of ${\mathfrak{M}}$ and Claims 3 and 4, every finite substructure ${\mathfrak{M}}$ satisfies (20), (21) and $\neg\mu$ and hence, by Claim 2, it embeds into ${\mathfrak{B}}$.
@@
Claim 6.
$\omega$ is pseudo cyclic and canonical with respect to the group $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Note that since ${\mathfrak{B}}$ is homogeneous in a finite relational signature, two $k$-tuples of elements of ${\mathfrak{B}}$ lie in the same orbit if and only if they satisfy the same atomic formulas. Therefore, the canonicity of $f$ and $g$ with respect to $\operatorname{Aut}({\mathfrak{B}})$ follows from the definition of ${\mathfrak{M}}$ and ${\mathfrak{N}}$: for $(a,b)\in B^{2}$, whether ${\mathfrak{B}}\models S(f(a,b))$ only depends on whether ${\mathfrak{M}}\models S(a,b)$ by Claim 5, which depends only on the atomic formulas that hold on $a$ and on $b$ in ${\mathfrak{B}}$. An analogous statement is true for atomic formulas of the form $R(x,y)$ and $x=y$. Therefore, $f$ is canonical. The argument for the canonicity of $g$ is analogous.
To see that $f$ and $g$ are pseudo cyclic, we show that $f^{*}$ and $g^{*}$ defined on $2$-orbits (using the terminology of Remark 7.4) are cyclic. By the definition of $f^{*}$, we need to show that for any $a_{1},a_{2},b_{1},b_{2}\in B$, the two pairs $(f(a_{1},b_{1}),f(a_{2},b_{2}))$ and $(f(b_{1},a_{1}),f(b_{2},a_{2}))$ satisfy the same atomic formulas. For the formulas of the form $S(x)$ and $R(x,y)$, this can be seen from Claim 5 and the definition of ${\mathfrak{M}}$ and ${\mathfrak{N}}$, since each of the conditions (22),(23),(24),(25),(21) and the union of (A), (B), (C), (D) is symmetric with respect to exchanging $x$ and $y$. For the atomic formulas of the form $x=y$, this follows from the injectivity of $f$. This shows that $f^{*}$ is cyclic; the argument for $g^{*}$ is the same. Hence, the pseudo-cyclicity of $f$ and $g$ is a consequence of Lemma 7.9 for $m=2$.
Claim 7.
$\omega$ improves $S$. By the definition of ${\mathfrak{M}}$ and ${\mathfrak{N}}$ and Claim 5, we have for all $x,y\in B$
@@
Claim 8.
$\omega$ improves $R$. Let $x_{1},y_{1},x_{2},y_{2}\in B$. We have to verify that
@@
  • ${\mathfrak{M}}\models\neg R((x_{1},y_{1}),(x_{2},y_{2}))$ and ${\mathfrak{N}}\models R((x_{1},y_{1}),(x_{2},y_{2}))$. By Claim 5, the left-hand side evaluates to $\frac{1}{2}$. By (22), we have ${\mathfrak{B}}\models\neg R(x_{1},x_{2})$ or ${\mathfrak{B}}\models\neg R(y_{1},y_{2})$. Therefore, the right-hand side of (26) is at least $\frac{1}{2}$ and the inequality holds.
@@
This exhausts all cases and concludes the proof of Claim 8.
It follows that $\omega$ is a binary fractional polymorphism of $\Gamma$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Polynomial-time tractability of $\operatorname{VCSP}(\Gamma)$ follows by Theorem 7.18 and 8.20. The final statement follows from Remark 8.21. ∎
@@ 9. Conclusion and Future Work
We have proved that if $\Gamma$ is a valued structure with an oligomorphic automorphism group and $R$ is a valued relation in the smallest valued relational clone that contains the valued relations of $\Gamma$, then $R$ is preserved by all fractional polymorphisms of $\Gamma$ (Lemma 6.9). We do not know whether the converse is true. It is known to hold for the special cases of finite-domain valued structures (Cohen et al., 2006; Fulla and Živný, 2016) and for classical relational structures with 00$\infty$ valued relations (CSP setting) having an oligomorphic automorphism group (Bodirsky and Nešetřil, 2003).
@@
In the following, we formulate a common generalisation of the complexity-theoretic implications of Conjecture 8.24 and the infinite-domain tractability conjecture from (Bodirsky et al., 2021) that concerns a full complexity classification of VCSPs for valued structures from reducts of finitely bounded homogeneous structures.
@@
Let $\Gamma$ be a valued structure with finite signature such that $\operatorname{Aut}(\Gamma)=\operatorname{Aut}({\mathfrak{B}})$ for some reduct ${\mathfrak{B}}$ of a countable finitely bounded homogeneous structure. If ($(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma$, then $\operatorname{VCSP}(\Gamma)$ is in P (otherwise, we already know that $\operatorname{VCSP}(\Gamma)$ is NP-complete by Theorem 3.4 and Corollary 5.14).
One might hope to prove this conjecture under the assumption of the infinite-domain tractability conjecture. Recall that also the finite-domain VCSP classification was first proven conditionally on the finite-domain tractability conjecture (Kolmogorov et al., 2017; Kozik and Ochremiak, 2015a), which was only confirmed later (Zhuk, 2017; Bulatov, 2017).
@@ Acknowledgements.
The first two authors have been funded by the European Research Council (Project POCOCOP, ERC Synergy Grant 101071674) and by the DFG (Project FinHom, Grant 467967530). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The third author was supported by the DFG project LU 1417/3-1 QTEC.
@@ Appendix A The Lebesgue Integral
It will be convenient to use an additional value $\infty$- ∞ that has the usual properties:
  • $-\infty<a$ for every $a\in\mathbb{R}\cup\{\infty\}$,
@@
The sum of $\infty$ and $\infty$- ∞ is undefined.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v6

@@ 8.6. An example of formerly open complexity
(recall that $((x_{1},y_{1}),(x_{2},y_{2}))$ might have been added to $R^{{\mathfrak{M}}}$ so that $\mathfrak{M}$} satisfies the sentence in (21)). The last option is false by the assumption and by (19), ${\mathfrak{B}}\models\neg S(x_{2})\wedge\neg S(y_{2})$, and hence neither (B) nor (D) holds. Therefore, one of the conditions (22), (A), or (C) holds for $((x_{1},y_{1}),(x_{2},y_{2}))$. Similarly, we obtain that one of the conditions (22) or (B) holds for $((x_{2},y_{2}),(x_{1},y_{1}))$, since ${\mathfrak{M}}\models R((x_{2},y_{2}),(x_{1},y_{1}))$ (to exclude (D) we use the assumption that ${\mathfrak{B}}\models S(x_{1})$ and hence $x_{1}\neq x_{2}$). This yields six cases and in each of them we must have that ${\mathfrak{B}}\models R(x_{1},x_{2})\wedge R(x_{2},x_{1})$ or ${\mathfrak{B}}\models S(y_{1})\wedge R(y_{1},y_{2})\wedge R(y_{2},y_{1})$. Since ${\mathfrak{B}}\models S(x_{1})\wedge R(x_{2},x_{2})\wedge R(y_{2},y_{2})$, this contradicts ${\mathfrak{B}}\models\neg\mu$. Since $(x_{1},y_{1}),(x_{2},y_{2})\in M$ were chosen arbitrarily, this shows that ${\mathfrak{M}}\models\neg\mu$. The argument for $\mathfrak{N}$} is similar. ∎
Claim 8.34.
There is an embedding $f$ of $\mathfrak{M}$} into $\mathfrak{B}$} and an embedding $g$ of $\mathfrak{N}$} into $\mathfrak{B}$}.
Proof.
We show the claim for $\mathfrak{M}$}; the proof for $\mathfrak{N}$} is analogous. By (Bodirsky, 2021, Lemma 4.1.7), it is enough to show that every finite substructure of $\mathfrak{M}$} embeds into $\mathfrak{B}$}. By the definition of $\mathfrak{M}$} and Claims  8.32 and 8.33, every finite substructure $\mathfrak{M}$} satisfies (20), (21) and $\neg\mu$ and hence, by Claim 8.31, it embeds into $\mathfrak{B}$}. ∎
@@
Claim 8.35.
$\omega$ is pseudo cyclic and canonical with respect to the group $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$.
Proof.
Note that since $\mathfrak{B}$} is homogeneous in a finite relational signature, two $k$-tuples of elements of $\mathfrak{B}$} lie in the same orbit if and only if they satisfy the same atomic formulas. Therefore, the canonicity of $f$ and $g$ with respect to $\operatorname{Aut}({\mathfrak{B}})$ follows from the definition of $\mathfrak{M}$} and $\mathfrak{N}$}: for $(a,b)\in B^{2}$, whether ${\mathfrak{B}}\models S(f(a,b))$ only depends on whether ${\mathfrak{M}}\models S(a,b)$ by Claim 8.34, which depends only on the atomic formulas that hold on $a$ and on $b$ in $\mathfrak{B}$}. An analogous statement is true for atomic formulas of the form $R(x,y)$ and $x=y$. Therefore, $f$ is canonical. The argument for the canonicity of $g$ is analogous.
To see that $f$ and $g$ are pseudo cyclic, we show that $f^{*}$ and $g^{*}$ defined on $2$-orbits (using the terminology of Remark 7.4) are cyclic. By the definition of $f^{*}$, we need to show that for any $a_{1},a_{2},b_{1},b_{2}\in B$, the two pairs $(f(a_{1},b_{1}),f(a_{2},b_{2}))$ and $(f(b_{1},a_{1}),f(b_{2},a_{2}))$ satisfy the same atomic formulas. For the formulas of the form $S(x)$ and $R(x,y)$, this can be seen from Claim 8.34 and the definition of $\mathfrak{M}$} and $\mathfrak{N}$}, since each of the conditions (22),(23),(24),(25),(21) and the union of (A), (B), (C), (D) is symmetric with respect to exchanging $x$ and $y$. For the atomic formulas of the form $x=y$, this follows from the injectivity of $f$. This shows that $f^{*}$ is cyclic; the argument for $g^{*}$ is the same. Hence, the pseudo-cyclicity of $f$ and $g$ is a consequence of Lemma 7.9 for $m=2$. ∎
Claim 8.36.
$\omega$ improves $S$.
Proof.
By the definition of $\mathfrak{M}$} and $\mathfrak{N}$} and Claim 8.34, we have for all $x,y\in B$
@@
Claim 8.37.
$\omega$ improves $R$.
Proof.
Let $x_{1},y_{1},x_{2},y_{2}\in B$. We have to verify that
@@
  • ${\mathfrak{M}}\models\neg R((x_{1},y_{1}),(x_{2},y_{2}))$ and ${\mathfrak{N}}\models R((x_{1},y_{1}),(x_{2},y_{2}))$. By Claim 8.34, the left-hand side evaluates to $\frac{1}{2}$. By (22), we have ${\mathfrak{B}}\models\neg R(x_{1},x_{2})$ or ${\mathfrak{B}}\models\neg R(y_{1},y_{2})$. Therefore, the right-hand side of (26) is at least $\frac{1}{2}$ and the inequality holds.
@@
This exhausts all cases and concludes the proof of Claim 8.37. ∎
It follows that $\omega$ is a binary fractional polymorphism of $\Gamma$ which is canonical and pseudo cyclic with respect to $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma)$. Polynomial-time tractability of $\operatorname{VCSP}(\Gamma)$ follows by Theorem 7.19 and 8.20. The final statement follows from Remark 8.22. This finishes the proof of Proposition 8.29.
@@ 9. Conclusion and Future Work
We have proved that if $\Gamma$ is a valued structure with an oligomorphic automorphism group and $R$ is a valued relation in the smallest valued relational clone that contains the valued relations of $\Gamma$, then $R$ is preserved by all fractional polymorphisms of $\Gamma$ (Lemma 6.9). We do not know whether the converse is true. It is known to hold for the special cases of finite-domain valued structures (Cohen et al., 2006; Fulla and Živný, 2016) and for classical relational structures with 0-$\infty$ valued relations (CSP setting) having an oligomorphic automorphism group (Bodirsky and Nešetřil, 2003).
@@
In the following, we formulate a common generalization of the complexity-theoretic implications of Conjecture 8.26 and the infinite-domain tractability conjecture (Bodirsky et al., 2021, Conjecture 1.2) that concerns a full complexity classification of VCSPs for valued structures from reducts of finitely bounded homogeneous structures.
@@
Let $\Gamma$ be a valued structure with finite signature such that $\operatorname{Aut}(\Gamma)=\operatorname{Aut}({\mathfrak{B}})$ for some reduct $\mathfrak{B}$} of a countable finitely bounded homogeneous structure. If $(\{0,1\};\operatorname{OIT})$ has no pp-construction in $\Gamma$, then $\operatorname{VCSP}(\Gamma)$ is in P (otherwise, we already know that $\operatorname{VCSP}(\Gamma)$ is NP-complete by Theorem 3.4 and Corollary 5.17).
One might hope to prove this conjecture under the assumption of the infinite-domain tractability conjecture. Recall that also the finite-domain VCSP classification was first proven conditionally on the finite-domain tractability conjecture (Kolmogorov et al., 2017; Kozik and Ochremiak, 2015a), which was only confirmed later (Zhuk, 2017; Bulatov, 2017). Note that Conjecture 9.3 does not generalize Conjecture 8.26, because it does not provide an algebraic condition for tractability. However, Conjecture 9.3 would imply a P vs. NP-complete dichotomy for resilience problems.
@@ Acknowledgements.
The first two authors have been funded by the European Research Council (Project POCOCOP, ERC Synergy Grant 101071674) and by the DFG (Project FinHom, Grant 467967530). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The third author was supported by the DFG project LU 1417/3-1 QTEC. This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/ESP6949724. The authors thank Antoine Mottet for pointing out a gap in the definition of composition of fractional maps, and Friedrich Martin Schneider for suggesting an elegant formalization of this composition.
@@ Appendix A The Lebesgue Integral
It will be convenient to use an additional value -$\infty$ that has the usual properties:
  • $\infty<a$ for every $a\in\mathbb{R}\cup\{\infty\}$,
@@
The sum of $\infty$ and -$\infty$ is undefined.
@@
$\int_{C^{D}}Xd\omega:=\sup\left\{\int_{C^{D}}Yd\omega\mid 0\leq Y\leq X,Y\text{ simple}\right\}.$
@@
Then both $X^{+}$ and $X^{-}$ are measurable, and both $\int_{C^{D}}X^{-}d\omega$ and $\int_{C^{D}}X^{+}d\omega$ take values in ${\mathbb{R}}_{\geq 0}\cup\{\infty\}$. If both take value $\infty$, then the integral is undefined (see Remark A.1). Otherwise, define
$\int_{C^{D}}Xd\omega:=\int_{C^{D}}X^{+}d\omega-\int_{C^{D}}X^{-}d\omega.$
@@
$E_{\omega}[X]$ $=\int_{C^{D}}X^{+}d\omega-\int_{C^{D}}X^{-}d\omega$
$=\sup\left\{\int_{C^{D}}Yd\omega\mid 0\leq Y\leq X^{+},Y\text{ simple}\right\}-\sup\left\{\int_{C^{D}}Yd\omega\mid 0\leq Y\leq X^{-},Y\text{ simple}\right\}$
@@
$\int_{{\mathscr{O}}^{(\ell)}_{C}}Xd\omega=\int_{{\mathscr{O}}^{(\ell)}_{C}}X^{+}d\omega-\int_{{\mathscr{O}}^{(\ell)}_{C}}X^{-}d\omega$
@@
$\int_{{\mathscr{O}}^{(\ell)}_{C}}X^{+}d\omega$ $=\sup\{\int_{{\mathscr{O}}^{(\ell)}_{C}}Yd\omega\mid 0\leq Y\leq X^{+},Y\text{ simple}\}$